OpenStudy (anonymous):

3/5 express as a decimal

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

\[ \frac{3}{5}*\frac{2}{2} = \frac{6}{10} = 0.6\] Say the fraction out loud: "six tenths". It's the same as the decimal 0.6

OpenStudy (anonymous):

thanks